# Diagnostic Reasoning in Natural Language: Computational Model and Application

### v1

@@ 6 Conclusion
We introduced NL-DAR as an important class of abductive reasoning problems and proposed a framework rooted in causality theory to approach NL-DAR problems. We outlined a practical guideline and toolkit to define and analyze NL-DAR problems. Our novel large-scale human dataset instantiates the framework for biomedical paper assessment. Our experiments showcase the versatility of our framework for analysis and LLM assistance. Our insights pave the path towards a broad study of natural language diagnosis tasks on the interface of causality, human-AI collaboration, and LLMs.
@@
Our framework for modeling NL-DAR problems is designed to be general and applicable to any natural language diagnostic reasoning task. We demonstrate its application and share the insights from our case study on biomedical paper assessment. More research is needed to validate the application of the NL-DAR framework for other diagnostic tasks.
At the core of our framework lies the definition of NL-DAR problems via a structural causal model. Here, we assume that the final verdicts arise when consuming the input text and do not exist on their own. This perspective is subject to philosophical discussion and most prominently advocated by Roland Barthes Barthes (1977). In practice, this means we assume that a text is consumed, causing the observed verdict (e.g., "the paper is a valuable contribution.") and that the reasoning steps in between give rise to this verdict. However, one might argue that there should be an objective verdict; for example, for paper assessment, a paper is either objectively correct or not. Likewise, one could argue that the verdict comes first, and the diagnosis is the post-hoc justification for it. While we acknowledge these points, we emphasize that the practical application of our framework in the case study supports our design choices in that diagnostic problems in natural language give rise to variance in the final verdict and in that the expert interviews support the directionality of causation that underlies our framework.
@@
The case study of this paper focuses on the biomedical domain. We believe that the developed SCM will be generally relevant to paper assessment in other scientific domains and can be used as a starting point for domain adaptation through expert interaction. Yet, the findings on the factors that shape human diagnosis are limited to the biomedical domain; researchers from other fields might put a different focus during the assessment; for example, the contribution of datasets and open resources is not considered in the biomedical paper assessment workflow, but is an important criterion in the assessment of NLP papers.
The data collection aims to cover a wide range of biomedical papers from various institutions and world regions to avoid biases toward a particular style of writing, scientific agenda, or scientific sub-community. However, some topical bias is inevitable to ensure relevance to the annotators’ expertise. We ensured high redundancy during annotation to avoid biasing the dataset to the idiosyncracies of individual annotators. Yet, the given sample of papers and the annotator pool might still introduce a certain level of bias to the dataset.
In the evaluation of LLMs as assistants for diagnostic reasoning, we rely on a range of automatic metrics for text generation. As our findings suggest, these metrics might not cover all aspects of a suitable answer in the workflow; for instance, they seem to favor more extensive answers, putting efficient but concise responses at a disadvantage. Likewise, the human baseline is a conservative estimate of true human performance; it is important to interpret our findings accordingly. Automatic metrics that consider the reasoning context of an answer, similar to the evaluation of chain-of-thought reasoning in LLMs Hao et al. (2024), are a promising direction of research to address these issues.
@@
This study has been approved by the university’s ethical board under the proposal number EK 41/2024. All senior annotators and experts participating in the study received adequate compensation within their regular employment in the research project. For the junior annotators, the workflow answers were collected as part of a university course aiming to teach students critical reading skills. The participation in the annotation study did not impact the students’ grading in the underlying course. All annotators could choose freely to reject participation in the study, and all human data was collected based on explicit consent.
@@
We explore the use of LLMs to automatically diagnose and evaluate biomedical papers. This technology has potential dual uses in peer review. A malicious party could use the LLM to skip the manual paper assessment process, which arguably requires the most effort during peer review. This is clearly problematic because LLMs cannot entirely substitute a human researcher, leading to incorrect or biased peer reviews. We stress that the output of such a system would be the answers to the reasoning steps and not a full peer-review report, which also covers questions to the authors or suggestions for improvement. We argue that the risk of misuse of this technology is relatively low. At the same time, the benefit of enabling new research on the reasoning underlying peer review and using paper assessment workflows as a teaching or assistance device for reviewers is considerable, and we believe that the benefits of the proposed technology outweigh the risks.
@@
Underlying our dataset is a set of publicly available papers with open licenses (CC-BY, CC0, …) that allow the redistribution and extension of the contents. However, as we publish subjective assessments of the papers by our study participants, this could potentially harm the reputation of the publication authors when they are assessed unfairly. We highlight that this is not the intended use of our dataset, which is exclusively meant for research on natural language diagnostic reasoning. We partially mitigate this risk by having at least three different researchers make annotations per paper: as a result, the collected dataset shows an opinion spectrum rather than a singular verdict on the paper. Hereby, we deem the risk of harming authors to lie below the day-to-day risk for reputational damage in regular scientific publishing; peer reviews are commonly released to the public, and papers are discussed (and criticized) publicly, for instance, on social media and in subsequent publications.
@@ Appendix A Appendix
A.1 Constructing an SCM for Biomedical Paper Assessment
A.1.1 Interviews and Formative Study

### v2

@@ 7 Conclusion
This paper introduced STRICTA– a novel specification framework for a family of structured reasoning problems over documents. We proposed a causality-based model, a practical guide, and a toolkit to help define and analyze such problems. Our novel, unique expert reasoning dataset instantiated the framework for biomedical paper assessment. Our experiments demonstrated the versatility of our framework for analysis and LLM assistance. Bridging the research in causality, human-AI collaboration, and LLMs, our study paves the path for more explainable, controlled, and grounded AI assistance for assessment tasks.
@@
Generalization and Domain Transfer
Our framework for STRICTA problems is designed to be general and applicable to any text assessment task. We demonstrate its application and share the insights from our case study on biomedical paper assessment. As peer review is one of the most challenging text assessment domains requiring deep expert reasoning and observing ambiguity and subjectivity, we deem it as most appropriate for the purpose of our study. We elaborate on the domain transfer and generality of STRICTA in Section 6.
@@
At the core of our framework lies the definition of STRICTA problems via a structural causal model. Here, we assume that the final verdicts and reasoning steps arise when consuming the input text and do not exist on their own. This perspective is subject to philosophical discussion Barthes (1977). In practice, this means we assume that a text is consumed, causing the verdict (e.g., "the paper is a valuable contribution.") and that the reasoning steps in between give rise to this verdict. However, one might argue that there should be an objective verdict; for example, for paper assessment, a paper is either objectively correct or not. While we acknowledge these points, we emphasize that the practical application of our framework in the case study supports our design choices in that text assessment problems give rise to variance in the final verdict and in that the expert interviews support the directionality of causation that underlies our framework.
@@
The case study of this paper focuses on the biomedical domain. The above transfer experiments in Section 6 show that the developed SCM is generally relevant to paper assessment in other domains. Yet, the findings on the factors that shape human text assessment are limited to the biomedical domain; researchers from other fields might put a different focus during the assessment; for example, the contribution of datasets and open resources is not considered in the biomedical paper assessment workflow, but is an important criterion in the assessment of NLP papers.
The data collection aims to cover a wide range of biomedical papers from various institutions and world regions to avoid biases toward a particular style of writing, scientific agenda, or scientific sub-community. However, some topical bias is inevitable to ensure relevance to the annotators’ expertise. We ensured high redundancy during annotation to avoid biasing the dataset to the idiosyncrasies of individual annotators. Yet, the given sample of papers and the annotator pool might still introduce a certain level of bias to the dataset. We highlight that the collected dataset is intended for research on reasoning during peer review across human subjects; it is not suitable for comparing behavioral patterns between groups, e.g. by varying levels of experience, which would require a different experimental design where annotators of different groups annotate the exact same papers.
LLM Assistance Experiments
In the evaluation of LLMs as assistants for structured reasoning, we rely on a range of automatic metrics for text generation. As our findings suggest, these metrics might not cover all aspects of a suitable answer in the workflow; for instance, they seem to favor more extensive answers, putting efficient but concise responses at a disadvantage. Likewise, the human baseline is a conservative estimate of true human performance; it is important to interpret our findings accordingly. Automatic metrics that consider the reasoning context of an answer, similar to the evaluation of chain-of-thought reasoning in LLMs Hao et al. (2024), are a promising direction of research to address these issues. We also conduct a human evaluation on a subset of the data to validate the automatic metrics. Structured reasoning offers built-in interpretability and facilitates human-AI collaboration by breaking down problems into smaller steps Luo et al. (2024); Ferdaus et al. (2024). The findings from our simulation study on human-AI collaboration support this principle. However, an important direction for future research is to investigate how LLMs and humans interact in real-world paper assessment scenarios. We perform a small-scale user study to validate that STRICTA-based assistance is practically useful during paper assessment (see A.2). However, to validate these findings, more and large-scale, dedicated user studies and performance metrics focused on human-computer interaction are needed, which lie beyond the scope of this work.
@@
This study has been approved by the university’s ethical board under the proposal number EK 41/2024. All senior annotators and experts participating in the study received adequate compensation within their regular employment in the research project. For the junior annotators, the workflow answers were collected as part of a university course aiming to teach students critical reading skills. The participation in the annotation study did not impact the students’ grading in the underlying course. All annotators could choose freely to reject participation in the study, and all human data were collected based on explicit consent. The annotators involved in the human evaluation received $10.40$£ per hour.
@@
We explore the use of LLMs to automatically assess biomedical papers. This technology has potential dual uses in peer review. A malicious party could use the LLM to skip the manual paper assessment process, which arguably requires the most effort during peer review. This is clearly problematic because LLMs cannot entirely substitute a human researcher, leading to incorrect or biased peer reviews. We stress that the output of such a system would be the answers to the reasoning steps and not a full peer-review report, which also covers questions to the authors or suggestions for improvement. We argue that the risk of misuse of this technology is relatively low. At the same time, the benefit of enabling new research on the reasoning underlying peer review and using paper assessment workflows as a teaching or assistance device for reviewers is considerable, and we believe that the benefits of the proposed technology outweigh the risks.
@@
Underlying our dataset is a set of publicly available papers with open licenses (CC-BY, CC0, …) that allow the redistribution and extension of the contents. However, as we publish subjective assessments of the papers by our study participants, this could potentially harm the reputation of the publication authors when they are assessed unfairly. We highlight that this is not the intended use of our dataset, which is exclusively meant for research on STRICTA. We partially mitigate this risk by having at least three different researchers make annotations per paper: as a result, the collected dataset shows an opinion spectrum rather than a singular verdict on the paper. Hereby, we deem the risk of harming authors to lie below the day-to-day risk for reputational damage in regular scientific publishing; peer reviews are commonly released to the public, and papers are discussed (and criticized) publicly, for instance, on social media and in subsequent publications.
Acknowledgments
This research work has been funded by the German Federal Ministry of Education and Research and the Hessian Ministry of Higher Education, Research, Science, and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE. It is co-funded by the European Union (ERC, InterText, 101054961). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This work benefited from the Federal Ministry of Education and Research (BMBF) projects “PlexPlain” (FKZ 01IS19081) and “XEI” (FKZ 01IS24079B). Further, we gratefully acknowledge the support of Microsoft with a grant for access to the OpenAI GPT models via the Azure cloud (Accelerate Foundation Model Academic Research). Finally, we express our gratitude to the members of the Synthetic RNA Biology Lab for their feedback, the numerous discussions, and their efforts during data annotation.
@@ Appendix A Appendix
A.1 Broader Related Work
We propose a framework for describing, collecting data, and assisting with structured reasoning during text assessment. Our study relates to a range of application domains of document quality assessment and general work on decomposed explicit human reasoning. Here, we extend the Related Work discussion in Section 2 to cover the broader related literature.
Explanations and Abductive Reasoning
We frame the LLM experiments on the dataset as an abductive reasoning task, aiming to discover the structured explanation for a given verdict. We contribute to the nascent area of abductive reasoning in NLP Bhagavatula et al. (2020). While prior work primarily focuses on abduction based on formal logic (e.g. Young et al., 2022), we focus on subjective reasoning on a real-world task. Relatedly, Dalal et al. (2024) evaluate LLMs’ ability to infer the best explanation for an observation and provide its reasoning as a free text. In our work, explanations consist of answers to steps in a reasoning graph imposing a consistent explanation structure. Our work relates to explainability research in NLP (e.g. Wiegreffe and Marasovic, 2021; Camburu et al., 2018). Notably, Dalvi et al. (2021) evaluate models on constructing reasoning paths (entailment trees) to explain science facts. From an explainability perspective, our framework generalizes entailment trees into directed graphs that provide more flexible and comprehensive explanations.
Diagnostic Reasoning
For this study we take inspiration in work on diagnostic reasoning Reiter (1987); Peng and Reggia (1990). Structured text assessment can be considered as a form of diagnostic reasoning where a given document is scrutinized to identify the underlying issues in the text that lead to a negative verdict on its quality, akin to a doctor’s diagnosis of a patient.
Previous studies in this domain focus on medical Richens and Buchard (2022); Magnani (2023) and engineering Koitz-Hristov and Wotawa (2018) problems, primarily using bipartite diagnosis graphs with numerical or categorical variables. In contrast, our framework considers multi-layered graphs, incorporates natural language variables, and leverages LLMs to model relationships between them.
Recently, clinical NLP research has increasingly focused on generating explanations for diagnoses to explicitly evaluate a model’s reasoning, which is most related to our study. Goh et al. (2024) study human-LLM collaboration in medical diagnosis, investigating whether clinicians’ diagnostic accuracy improves when supported by ChatGPT. In our work, we simulate human-LLM collaboration during diagnosis but limit LLM support to individual reasoning steps. This approach provides more granular insights into the utility of LLMs for various aspects of diagnostic reasoning. Savage et al. (2024) use LLMs in a zero-shot setting, designing prompts to mimic medical expert reasoning during diagnosis. Complementing this, our study does not focus on prompt design but instead introduces an LLM program architecture informed by expert reasoning processes. Most related to our work are prior studies that use explicit reasoning structures to guide LLMs. For instance, Yang et al. (2024) develop DrHouse, a tool for interactive anamnesis and diagnosis using LLMs augmented with a medical knowledge base. They structure LLM prompts with diagnosis guidelines as binary decision trees to support follow-up questioning and eventual diagnosis. Similarly, Wang et al. (2024a) propose an augmented LLM-program architecture that iteratively generates a chain of diagnostic steps, forming a reasoning trace for the final diagnosis in the style of entailment trees Dalvi et al. (2021). In contrast, our study employs directed acyclic graphs to represent reasoning structures. Unlike trees, these graphs allow reasoning steps to recombine multiple prior observations, offering a more flexible framework. Overall, we complement the field of clinical diagnostic reasoning by applying similar techniques to a new domain – diagnosing texts instead of patients – and propose a modeling framework that accommodates the subjectivity and ambiguity inherent to the textual domain, which differs from the medical diagnosis context.
Essay Grading
Automatic essay grading aims to score student essays’ quality and is usually cast as a regression of classification problem Misgna et al. (2025). Most related to our study are approaches on feedback generation serving as an actionable explanation for the verdict and that score along a rubric Wang et al. (2024b). For example, Han et al. (2023) score essays along multiple dimensions serving as an explanation for the overall quality assessment, but they do not consider reasoning structures. Wang et al. (2024b) design a human-LLM collaboration system for rubric-based scoring, but they do not structure the process. More structured, Fiacco et al. (2022) design a system to annotate essays with rhetorical tree structures serving as a means to identify problematic areas in the essay and produce a quality judgement, but they do not consider complex reasoning over these structures. Our framework connects the paper to the output, so it comes with more expressive explanations and is more fine-grained.
A.2 Discussion: Paper Assessment as a Explicit Structured Reasoning Process
This Section offers an in-depth discussion of the newly proposed perspective on document and paper assessment as a structured reasoning task.
Paper Assessment as a STRICTA problem
We frame paper assessment as a structured reasoning problem in critical text assessment. This framing is plausible and flexible with respect to the partially subjective nature of paper assessment. Specifically, we motivate this framing from the expert discussions and from an analysis of peer review data.
The expert interviews and formative study (see A.3.1) revealed that assessing a biomedical paper is largely independent of its specific topic. Experts repeatedly investigate the same components, considering the line of argument and underlying research of the paper to determine its quality. Each review differs only in the weight that is assigned to the different aspects, e.g., leaving some out sometimes when irrelevant to the paper type, and how each aspect is determined. This justifies framing paper assessment from the structured reasoning angle because an interconnected set of reasoning variables adequately describes the experts’ model of paper quality.
Besides the evidence derived from the expert interviews, we validate the proposed framing within the context of peer review. We analyze $30$ randomly sampled reviews from the NLP domain of the NLPEER corpus Dycke et al. (2023) and manually analyze their lines of argument. This analysis reveals that reviewers often employ what-if-reasoning (e.g. "Without additional experiments covering more models, the expressiveness of the experiments is limited." implying that if the paper had more experiments the reviewer would value the soundness of the paper differently) and that many weaknesses identified by reviewers reoccur across a variety of papers. These observations support the framing of paper assessment as a structured reasoning process and a strong alignment with the causal framework proposed in this paper.
Advantages of the STRICTA Framing
Framing paper assessment as a STRICTA problem opens up safe and interpretable human-AI collaboration. In such a scenario, the human expert reads the paper and defines selected variables, and the AI is tasked to fill in the remaining variables automatically or in collaboration with the human. In our experiments, we test different configurations, including providing only the final verdict, and providing full agency to the human on the outcome decision. Since the reasoning structure for the explanation is fixed, the model output becomes easier to interpret.
Relation to Other Forms of Reasoning
The STRICTA framework allows for defining various reasoning tasks on the underlying workflow depending on the provided data. In the following, we discuss the reasoning task considered during the experiments in Section 5 – the task of explaining a given verdict based on the paper text – in reference to the broader reasoning concepts.
The work is inspired by diagnostic reasoning. Diagnosis is a general activity that is not limited to medical diagnosis by physicians. More generally, diagnosis refers to the task of identifying the hidden cause of an observed (malfunctioning) behavior in a system Reiter (1987). For instance, psychotherapists, programmers during debugging of source code, or engineers analyzing machine failures all use diagnostic reasoning (e.g. Mueller et al., 2012; Wong et al., 2016; Hameed et al., 2009; Magnani, 2023). By the above general definition, paper assessment can be seen as a diagnostic task. Figure 4 illustrates this in comparison with other diagnostic activities.
Finally, we provide a short discussion on the classification of the considered reasoning task. Abductive reasoning means inferring the most likely explanation for a given set of observations Peirce (1955). Deductive reasoning means deriving a conclusion from a set of premises by applying a sequence of logic rules Leighton (2003). We argue that the given problem setting is abductive reasoning rather than deductive. First, unlike deduction, the reasoning process of paper assessment is probabilistic, i.e., reviewers choose the best of all possible explanations for their verdict. The verdict is not derived from a set of formal rules from a set of premises. The subjective nature of the task and the complex interaction of quality criteria make this impossible. Second, in the given task, framing the goal is to find the unobserved intermediate steps that lead to the observed verdict, which aligns precisely with the definition of abduction.
Figure 4: Diagnosis is a general concept used by many professions. Scientists use it during paper assessment (top), physicians use it during medical diagnosis (middle), and engineers apply it to analyze machine failures (bottom). An observed failure behavior (yellow) given the observable system states (red) is explained by assigning the hidden system states a most likely value (grey area).
Practical Utility of STRICTA-based Assistance
Conceptually, STRICTA supports effective and interpretable human–AI collaboration in two ways: by enabling explainability and identifying opportunities for human–LLM collaboration. Structured explanations enhance interpretability and explainability of NLP models by augmenting single-shot black-box predictions with intermediate reasoning steps, such as reasoning trees Dalvi et al. (2021) or multi-hop predictions Jiang et al. (2020). By design, they allow verification of model outputs, promoting trustworthiness. Structured reasoning also enables systematic analysis of human–LLM collaboration potential (see Section 5). It helps identify steps needing assistance, assess LLM competence across reasoning stages, and estimate the feasibility of dynamic task routing between humans and AI.
Beyond these conceptual contributions, we empirically test STRICTA’s practical utility during paper assessment. We conduct a small-scale user study with nine postgraduate biomedical students who assess preprints in their field using the dataset generation interface, augmented with GPT-3.5 assistance. Participants could freely engage with the workflow and request AI support at any step. Six completed a post-study survey. On a scale from 1 (fully disagree) to 7 (fully agree), the structured workflow received an average rating of 5.67 for usefulness in paper assessment. The utility of AI assistance during stepwise evaluation was rated 3.8 on a scale from 1 (worst) to 5 (best). These results indicate the potential of STRICTA for effective and satisfactory human–LLM collaboration in text assessment. However, more analysis on the human-LLM interaction dynamics based on large-scale controlled user studies are needed to validate the results of this small-scale study.
A.2.1 Background on Causality
Following the definition by Pearl (2000):
Definition 2
An SCM $\mathcal{M}$ is $(U,V,F,P_{\mathcal{M}})$, where
  1. i.
    $U$ is a set of exogenous variables determined by factors outside of $\mathcal{M}$. $P_{\mathcal{M}}$ is a probability function over the domain of $U$.
  2. ii.
  3. iii.
    $F=\{f_{1},...,f_{n}\}$ is a set of functions, the structural equations, on the domains of $U_{i}\cup\textit{pa}_{i}$ to $V_{i}$, where $\textit{pa}_{i}\subseteq V\setminus V_{i}$, and $U_{i}\subseteq U$ and $F$ forms a mapping from $U$ to $V$.
Causal Graphs
Interventions and Counterfactuals
We notate interventions using the $do$-operator. For an intervention $do(X=x)$, we alter the SCM $\mathcal{M}$ by replacing all occurrences of variable $X$ by $x$ in $F$. We denote the altered model as $\mathcal{M}_{X=x}$. Counterfactuals pose what-if questions on the SCM. For a query "What value had $Y$ if $X=x$ under the circumstances $u$?" we imagine the causal effect of $X$ on $Y$ for a concrete instance counter to the observed facts. The counterfactual distribution is
$P_{\mathcal{M}_{X=x}}(Y=y)=\sum_{\{u|u\in\widehat{U}\land Y_{x}(u)=y\}}P(u)$
where $Y_{x}(u)$ denotes the potential response of $Y$ under $\mathcal{M}_{X=x}$. Finally, the average causal effect (ACE) of an intervention on a binary variable $X$ is the difference of expected values of $Y$ with $\textit{ACE}(X,Y)=\mathbb{E}[Y_{x=1}]-\mathbb{E}[Y_{x=0}]$.
A.2.2 STRICTA Design Choices
Characterizing Reasoning Problems
For the discussion of alternative formalisms for describing STRICTA  problems we refer to the literature on the nascent field of diagnostic abductive reasoning (DAR) Peng and Reggia (1990) since it also aims to make reasoning explicit and expresses the reasoning as a structured graph.
One very established approach is to model DAR as a set-covering problem Bylander et al. (1991). Here, the task is to select a minimal set of hypotheses that explain the observed behavior optimizing for a plausibility metric. Under this framing, the system model consists of a set of independent boolean hypotheses and their relation to the observable behavior. As another formalism, propositional logic has been proposed to describe the system model Nordh and Zanuttini (2008). Here, the hypothesis space is expressed as a set of propositional clauses from which a subset needs to be selected such that it is faithful to the problem context and logically implies the observations. The system model hereby consists of multiple propositional logical clauses and solving the task means finding a satisfiable configuration of hypotheses given the observations. Similarly, logic programming models the system as a logic program, posing the behavioral variables as known variables and searching for an optimal answer Lin and You (2002).
While the above approaches model reasoning as a purely logical and thereby discrete problem, another line of research focuses on probabilistic models of abductive reasoning that explicitly incorporate ambiguity and uncertainty and often support continuous variables. Fuzzy logic models the space of explanations in terms of fuzzy sets, and the solution is discovered by maximizing the probability of the hypothesis set explaining the observable behavior Berner (2007); Suryanarayanan et al. (1995). More prominently, Bayesian networks and structural causal models Richens et al. (2020). Here, the underlying causes for observable behavior are expressed in a network of interacting factors. While all formalisms allow for the encoding of a system of components that jointly influence behavioral variables, logic-based approaches cannot handle ambiguity and uncertainty, whereas statistical approaches can only characterize associational strengths between factors. Only SCMs allow for studying cause and effect relationships Pearl (2000), making them strictly more powerful than the other approaches.
Design Choices
The key challenge in approaching a concrete STRICTA problem or family of STRICTA problems lies in choosing an appropriate structured reasoning model (i.e., the workflow) which is sufficiently expressive yet computational inferences on this model are feasible.
We opt to describe STRICTA problems using SCMs. On one hand, the probabilistic elements of this formalism can directly model the statistical variance induced by the ambiguity and subjectivity of most document assessment tasks. On the other hand, the ability to pose causal and counterfactual queries on SCMs enables a rich methodology to study the reasoning process of humans and potentially AI models in terms of explanations.
Additionally, for any reasonable text assessment process, the final verdict should depend on the input text of the problem. Therefore, we require the SCM for an STRICTA problem to have three tiers, each depending on the previous tier: the surface text variables, the intermediate reasoning steps, and the final verdict.
A.2.3 STRICTA Beyond Paper Assessment
To exemplify the transfer of STRICTA  to other tasks, we outline the steps necessary to instantiate it for multi-rubric essay scoring. Automatic essay scoring (AES) can enhance teacher consistency and enable large-scale student self-teaching Taghipour and Ng (2016). Beyond predicting an overall score, AES models often provide rubric-specific scores that relate to the final grade Kumar et al. (2022). This task aligns well with STRICTA’s scope, and we propose a streamlined protocol based on our study. To instantiate STRICTA for essay scoring, we follow these steps (see Section 3.3): (1) design the causal graph, (2) populate it with human data, and (3) perform analysis and partial automation.
To construct the causal graph (1), we first identify the relevant variables, starting with the final verdict as the final score along with a rationale. We model the rubric scores as infer steps that are direct parents to the final verdict. The key challenge lies in defining the information flow from the essay text via extract steps towards the scoring rubrics. Here, we may involve teachers and grading experts to understand their mental workflow. As an example, a rubric score on “narrative quality” Somasundaran et al. (2018) is determined by first extracting the exact narrative of the essay and then deriving partial judgements on its consistency and alignment with the essay prompt. As an alternative to the manual definition of the workflow, existing resources of rubric assessments (e.g. Yoo et al. (2024)) can be used to determine the different factors impacting essay scoring and their relation to the text. To populate the structural causal model with data (2), we replicate the annotation study with human experts based on existing essay corpora, tasking teachers to follow the workflow to grade an essay. Based on our findings, a human-LLM collaborative approach to data generation is promising to speed up this process, and a combination of authentic human data as well as LLM-generated data revised by humans is promising. Finally, we perform an analysis and evaluate LLMs on the task (3). The analysis can be used to quantify the consistency of human ratings (through variability analysis), determine the weighing of rubrics and their localization in the essay text (through average causal effect analysis). Applying LLMs for automatic scoring according to the grading workflow enables students to inspect the specific rationales for the final score and to locate them in the paper. Conceivably, students can use the resulting tool to iteratively improve their paper by producing counterfactual versions that improve their score as a learning exercise.
Overall, this example illustrates the application of STRICTA  to other critical text assessment scenarios and makes clear the potential for analysis and automation in this domain.
A.3 Constructing an SCM for Biomedical Paper Assessment
A.3.1 Interviews and Formative Study